# FML-bench: A Benchmark for Automatic ML Research Agents Highlighting the Importance of Exploration Breadth

### v1

@@ Appendix B Task Settings
We adopt the AIF360 (Bellamy et al., 2018) repository with Adversarial Debiasing as the baseline, using the Adult (Becker & Kohavi, 1996) dataset. The primary fairness metric is average odds difference, and the agent is tasked with enhancing fairness (minimizing absolute average odds difference) while maintaining or improving classification accuracy. This task measures the agent’s ability to balance equitable outcomes with model utility across demographic groups.
Appendix C Evaluation Metrics
Utility reflects empirical improvement within each task and is reported using the task-specific performance metric.
Diversity quantifies implementation dispersion within the best-performing round. Given $i$ th step code embedding $\mathbf{e}_{i}$ extracted from CodeGraphBERT (Guo et al., 2020) and centroid $\bar{\mathbf{e}}=\frac{1}{n}\sum_{i=1}^{n}\mathbf{e}_{i}$, diversity is computed as $\frac{1}{n}\sum_{i=1}^{n}\|\mathbf{e}_{i}-\bar{\mathbf{e}}\|_{2}$. Greater dispersion indicates broader exploration of implementation choices within a single round.
Academic Contribution Rate reflects the agent’s tendency toward innovative algorithmic contributions versus conventional engineering optimizations. It is defined as $N_{\text{aca}}/N_{\text{suc}}$, where $N_{\text{aca}}$ is the number of steps recogniezd by Qwen3 as academic modifications (e.g., new losses or architectures) relative to the baseline, and $N_{\text{suc}}$ is the number of steps whose experiments execute without errors and yield valid results.
Step Success Rate is defined as the $N_{\text{suc}}/N_{\text{comp}}$. Combining the metric Step Completion Rate ($N_{\text{comp}}/N_{\text{total}}$) will reflect the system reliability of the agent. Here, $N_{\text{comp}}$ is the number of steps actually executed and $N_{\text{total}}$ is the number of steps assigned to the agent.
Cost is reported as total token consumption (sum across steps) and step token consumption (average per step), together with total time consumption and step time consumption, to characterize computational and temporal efficiency.
Step To Target calculation Steps To Target is measured as $STT=\min\{n:\text{performance}(n)\geq\text{threshold}\}$, representing the minimum number of iterations required to reach the target performance level, where the threshold is set to be the worst best-run performance among the agents.
C.1 Academic modification Classification
Academic Contribution Rate is computed by identifying academic steps within the best-performing rounds. We use Qwen3 (version: 235b-a22b-2507) to automatically classify each modification as academic or engineering by comparing the baseline and modified code. The prompt used to guide Qwen3 is provided below.
You are an experienced machine learning researcher. Your task is to analyze two pieces of code—baseline and modified—and judge whether the modification is mostly prone to engineering or academic modification. You will also assign one subtag to describe the specific type(s) of modification. Baseline code is delimited by [baseline:begin] …[baseline:end]. Modified code is delimited by [modified:begin] … [modified:end]. Carefully compare the two kinds of code and reason about what changed and why it matters. Please give your judgement by ’engineering’ or ’academic’ and one the most related subtag eg: [engineering, [ENG/LR_SCHED]], [academic, [ACD/LOSS_NEW]]. Your answer should only select one of the two options [engineering] or [academic]. Engineering modification: Changes focused on making an existing method run more stably, efficiently, or accurately through implementation details, system-level tuning, or configuration, without introducing a new learning principle, objective family, or architectural paradigm. Typical signals: hyperparameter tuning, training schedules, efficiency and stability tricks, data pipeline and evaluation scripting that keep the task/method essentially the same. Academic modification: Changes that introduce or test a new machine learning idea: a new objective/loss with motivation, new inductive bias or architecture, a different training or inference paradigm, etc. For Engineering subtags: [ENG/LR_SCHED]: Learning-rate, momentum, scheduler (Cosine, OneCycle, Warmup) [ENG/OPT_SWAP]: Optimizer swap as tuning (SGD to AdamW) without new update rule. [ENG/LOSS_WEIGHTING]: Loss weights/temperature/threshold sweeps (no new loss family). [ENG/TRAIN_STR]: Early stopping or extend training time. [ENG/DATA_AUG]: Data cleaning/dedup/resampling; standard/strong augmentation parameters For Academic subtags: [ACD/LOSS_NEW]: New loss/constraint; dual/information-theoretic objectives [ACD/ARCH_NEW]: New layer/module/architecture (e.g., novel attention; new encoder–decoder coupling) [ACD/TRAIN_PARADIGM]: New training or inference paradigm (e.g., contrastive to masked/generative; new decoding strategy; beyond teacher forcing; supervised to self-supervised) Your answer should only select one of the two options [engineering] or [academic] and one the most related subtag eg: [engineering, [ENG/LR_SCHED]], [academic, [ACD/LOSS_NEW]]. [baseline:begin] … baseline code … [baseline:end] [modified:begin] … modified code … [modified:end] Compare baseline code and modified code, tell me which kind of code modification it is to improve the baseline method, choose from engineering, academic and one most related subtag eg: [engineering, [ENG/LR_SCHED]], [academic, [ACD/LOSS_NEW]]
Appendix D Protecting Evaluation Integrity
@@
Appendix E Additional Results
@@
This section provides a comprehensive comparative analysis of automated AI research systems, focusing on both their performance and operational efficiency. We evaluate TheAIScientist, AIDE, and Claude Code across eight core machine learning research problems, considering not only their final outcomes but also their operational behavior over time. Fig. 3 depicts the step-wise performance progression of these systems across all tasks, while Fig. 4 highlights the relationship between performance and solution diversity.
Our analysis covers multiple key dimensions: solution diversity (Tab. 4), research contribution quality (Tab. 5), computational efficiency in terms of total token usage (Tab. 6) and step-wise token usage (Tab. 7), runtime efficiency in terms of total time (Tab. 8) and step-wise time (Tab. 9), operational reliability measured by step success rate (Tab. 10) and step completion rate (Tab. 11), and improvement speed (Tab. 12). Together, these results provide detailed insights into the trade-offs between different agentic research designs, illustrating how they influence both the quality of research outcomes and the efficiency of resource utilization.
Figure 4: Performance - Diversity Analysis.
Table 4: Code Diversity Comparison of Best Performance Round. G2.5-Pro denotes Gemini-2.5-Pro. Bold numbers indicate the best (highest) performance in each row.
@@
Generalization 46.38 17.94 0.00 21.73 22.25
Data Eff. 25.43 17.70 18.33 11.06 17.98
Rep. Learn. 17.42 10.86 15.06 8.76 4.74
Cont. Learn. 21.41 45.23 30.79 18.03 5.42
Causality 19.22 18.25 17.51 8.19 7.06
Robust. & Rel. 16.65 30.13 29.88 56.05 25.90
Privacy 33.84 13.13 25.29 10.13 5.93
Fair. & Bias 19.29 12.06 25.23 13.98 6.91
Average $\pm$ Std 24.95±9.63 20.66±10.85 20.26±9.37 18.49±14.87 12.02±8.04
Table 5: Academic Contribution Rate (ACR) Comparison. G2.5-Pro denotes Gemini-2.5-Pro. Bold numbers indicate the best (highest) performance in each row.
@@
Generalization 0.90 0.95 0.33 0.63 0.33
Data Eff. 1.00 0.89 0.97 0.99 1.00
Rep. Learn. 0.93 0.93 0.98 0.99 0.00
Cont. Learn. 0.12 0.17 0.89 0.35 0.00
Causality 0.88 0.78 0.93 0.89 0.00
Robust. & Rel. 0.95 0.87 0.92 0.33 0.67
Privacy 0.90 0.83 0.80 0.07 0.00
Fair. & Bias 0.95 0.84 0.93 0.95 0.00
Average $\pm$ Std 0.83±0.27 0.78±0.24 0.84±0.20 0.65±0.34 0.25±0.36
@@
Generalization 9.34 9.88 0.18 3.80 5.98
Data Eff. 3.09 2.55 0.33 1.53 7.98
Rep. Learn. 4.69 3.99 0.40 1.51 8.89
Cont. Learn. 7.81 7.74 1.06 3.21 2.99
Causality 4.57 3.05 0.56 2.99 9.26
Robust. & Rel. 7.24 6.74 1.69 2.93 8.22
Privacy 6.64 4.45 1.68 1.61 13.88
Fair. & Bias 5.13 4.21 0.94 3.12 9.34
Average $\pm$ Std 6.07±2.05 5.33±2.55 0.85±0.59 2.59±0.90 8.32±3.10
@@
ML Problems TheAIScientist AIDE Claude Code
GPT-5 G2.5-Pro GPT-5 G2.5-Pro Opus-4.1
Generalization 0.09 0.10 0.09 0.10 1.99
Data Eff. 0.03 0.03 0.01 0.02 0.50
Rep. Learn. 0.05 0.04 0.01 0.01 0.81
Cont. Learn. 0.08 0.08 0.04 0.05 1.49
Causality 0.05 0.03 0.02 0.03 0.77
Robust. & Rel. 0.07 0.07 0.03 0.05 1.64
Privacy 0.07 0.04 0.02 0.02 4.63
Fair. & Bias 0.05 0.04 0.02 0.03 2.34
Average $\pm$ Std 0.06±0.02 0.05±0.03 0.03±0.03 0.04±0.03 1.77±1.32
@@
Fair. & Bias 7.05 3.46 1.32 4.46 0.25
Average $\pm$ Std 8.32±3.15 12.23±7.93 2.46±1.77 8.06±6.44 1.27±1.77
@@
Fair. & Bias 4.23 2.07 2.02 2.63 3.80
Average $\pm$ Std 10.39±10.79 12.73±14.64 10.79±12.86 8.65±7.92 12.49±10.77
ML Problems TheAIScientist AIDE Claude Code
GPT-5 G2.5-Pro GPT-5 G2.5-Pro Opus-4.1
Generalization 0.7700 0.9900 0.5000 0.7027 1.0000
Data Eff. 0.9600 0.6500 0.2333 0.2871 0.9375
Rep. Learn. 0.8300 0.8400 0.8000 0.6040 0.5455
Cont. Learn. 0.9600 0.9400 0.8214 0.9155 1.0000
Causality 0.8900 0.9000 0.8214 0.9406 0.8333
Robust. & Rel. 0.9900 0.9500 0.8281 0.6667 0.6000
Privacy 0.9700 0.9500 0.8714 0.9571 1.0000
Fair. & Bias 0.9500 0.7200 0.3590 0.7426 0.7500
Average $\pm$ Std 0.92±0.08 0.87±0.12 0.65±0.25 0.73±0.22 0.83±0.18
ML Problems TheAIScientist AIDE Claude Code
GPT-5 G2.5-Pro GPT-5 G2.5-Pro Opus-4.1
Generalization 1 1 0.02 0.36 0.03
Data Eff. 1 1 0.30 1 0.16
Rep. Learn. 1 1 0.45 1 0.11
Cont. Learn. 1 1 0.28 0.70 0.02
Causality 1 1 0.28 1 0.12
Robust. & Rel. 1 1 0.64 0.53 0.05
Privacy 1 1 0.70 0.69 0.03
Fair. & Bias 1 1 0.39 1 0.04
Average $\pm$ Std 1.00±0.00 1.00±0.00 0.38±0.22 0.79±0.25 0.07±0.05
Table 12: Improvement Speed Comparison. We report the number of steps used to reach the threshold, which is defined as the best result achieved in the best run of the worst-performing agent. Fewer steps indicate higher improvement speed. G2.5-Pro denotes Gemini-2.5-Pro. Bold numbers indicate the best (lowest) performance in each row.
ML Problems TheAIScientist AIDE Claude Code
GPT-5 G2.5-Pro GPT-5 G2.5-Pro Opus-4.1
Generalization 1 1 1 1 1
Data Eff. 1 2 1 1 3
Rep. Learn. 29 10 15 40 10
Cont. Learn. 6 1 1 1 2
Causality 13 38 17 7 6
Robust. & Rel. 26 50 5 34 2
Privacy 47 83 31 40 2
Fair. & Bias 1 25 14 17 4
Average $\pm$ Std 15.50±16.95 26.25±29.40 10.63±10.66 17.63±17.78 3.75±2.96
Appendix F Prompts Details
This section presents the detailed prompt specifications that form the foundation of our autonomous research agent framework. The prompts serve as the primary interface between human researchers and AI agents, translating high-level research objectives into actionable instructions that can guide systematic scientific inquiry across diverse machine learning domains. The prompt design philosophy centers on creating a structured yet flexible research environment that balances autonomy with scientific rigor. Rather than providing overly prescriptive instructions that limit creative exploration, these prompts establish clear boundaries, evaluation criteria, and operational constraints while encouraging the agent to develop and test novel hypotheses within established research paradigms.
@@ Appendix F Prompts Details
Second, we detail the Autonomous Research Agent Framework that governs how agents interact with research codebases to conduct iterative experimentation. This operational framework transforms the conceptual research challenges into executable workflows, ensuring that agent behavior follows sound scientific methodology while maintaining reproducibility and experimental integrity. Together, these prompt specifications create a comprehensive research environment where autonomous agents can contribute meaningfully to advancing machine learning across multiple disciplines, providing both the research contexts and the methodological framework necessary for systematic scientific progress.
@@
The following task descriptions establish comprehensive research contexts spanning 8 fundamental areas of machine learning. Each prompt follows a structured format that defines: (1) the researcher’s identity and expertise, (2) the specific technical setup, including datasets and baseline methods, (3) clear optimization objectives and constraints, and (4) fairness criteria to ensure meaningful comparisons.
These prompts span a wide range of machine learning challenges, including generalization, data efficiency, privacy, fairness, and robustness. Each task is anchored in established benchmarks and frameworks, such as DomainBed for domain generalization, EasyFSL for few-shot learning, and AIF360 for fairness-aware learning, providing realistic experimental environments that closely mirror actual research workflows.
The prompts are carefully crafted to encourage both incremental improvements to existing methods and the exploration of novel algorithmic approaches, while maintaining scientific rigor through controlled experimental conditions. This design enables systematic evaluation of how autonomous agents can contribute to advancing the state-of-the-art across multiple ML disciplines simultaneously.
@@
System: You are an ambitious AI PhD student focused on advancing machine learning for causal inference, reasoning, and interpretable modeling. Task Description: You are working with the Dragonnet framework to estimate individual treatment effects (ITEs) in both real (IHDP) and simulated data scenarios. The simulated data follows Setup A from Nie & Wager (2018), featuring difficult nuisance functions (e.g., propensity scores) but simple, easily identifiable treatment effects. Your goal is to improve the precision of treatment effect estimation across both IHDP and simulated benchmarks.
@@
System: You are an ambitious AI PhD student focused on improving fairness-aware learning with AIF360’s Adversarial Debiasing on the Adult dataset. Task Description: You are working with AIF360’s Adversarial Debiasing (classifier–adversary) as the baseline on the Adult dataset to evaluate the fairness–accuracy trade-off. Your goal is to minimize absolute Average Odds Difference toward parity (=0) while maintaining or improving Balanced Accuracy on held-out test splits and across protected subgroups (e.g., sex/race). You should enhance the baseline Adversarial Debiasing algorithm, but you may also propose entirely new fairness methods if they better support reduced absolute Average Odds Difference without sacrificing Balanced Accuracy. You are allowed to refine the classifier and adversary networks and the training pipeline, provided comparisons remain fair to the original setup (similar capacity, training budget, and data access). The priority is minimizing absolute Average Odds Difference while preserving or improving Balanced Accuracy.
Prompting Claude Code as an Autonomous Research Agent

### v2

@@ Appendix A Task Design Details
This task measures the ability to balance equitable outcomes across demographic groups with overall model utility in classification settings. We selected the COMPAS dataset (Angwin et al., 2016), a recidivism prediction benchmark that presents meaningful fairness challenges due to documented disparities across protected attributes such as race and sex. As the baseline, we adopted Adversarial Debiasing, which employs a classifier-adversary architecture to learn fair representations by preventing the adversary from predicting sensitive attributes. We constructed this task using the AIF360 repository (Bellamy et al., 2018). The evaluation metric is absolute average odds difference, with the objective of minimizing this disparity toward parity (zero) while maintaining or improving classification accuracy across protected subgroups.
Appendix B Case Analysis
B.1 Worst Case Analysis
Modification Distribution by Agents
As shown in Tab. 5, TheAIScientist strongly favors algorithmic modifications (56.2%) with zero implementation errors, indicating robust execution. AIDE presents a more balanced modification distribution but experiences higher implementation and execution issues (31.2%). Claude Code shows elevated parameter configuration rates (37.5%), equivalent to its algorithmic modification rate.
Modification Distribution by LLMs
As shown in Tab. 6, Gemini-2.5-Pro strongly favors algorithmic modifications (68.8%), preferring core algorithm changes; GPT-5 primarily focuses on structural modifications (50.0%) with emphasis on architecture adjustments; while Opus-4.1 maintains balanced distribution between parameter configuration and algorithmic modifications.
B.2 Best Case Analysis
Modification Distribution by Agents
As shown in Tab. 7, TheAIScientist shows strong algorithmic innovation focus, predominantly targeting core algorithmic logic; AIDE combines strong algorithmic innovation with hyperparameter tuning; while Claude Code uniquely favors architectural modifications, including layer additions and normalization adjustments.
Modification Distribution by LLMs
As shown in Tab. 8, GPT-5 exclusively prioritizes algorithmic modifications (100.0%); Gemini-2.5-Pro exhibits the most diversified approach, uniquely incorporating substantial data augmentation efforts (28.6%); while Opus-4.1 balances attention between architectural and algorithmic modifications.
Table 5: Worst Case Modification Distribution by Agents
Modification Type TheAIScientist AIDE Claude Code
Algorithmic Modification 56.2% (9) 37.5% (6) 37.5% (3)
Structural Modification 31.2% (5) 25.0% (4) 12.5% (1)
Parameter Configuration 12.5% (2) 6.2% (1) 37.5% (3)
Implementation & Execution 0.0% (0) 31.2% (5) 12.5% (1)
Table 6: Worst Case Modification Distribution by LLMs
Modification Type GPT-5 Gemini-2.5-pro Opus-4.1
Algorithmic Modification 25.0% (4) 68.8% (11) 37.5% (3)
Structural Modification 50.0% (8) 6.2% (1) 12.5% (1)
Parameter Configuration 6.2% (1) 12.5% (2) 37.5% (3)
Implementation & Execution 18.8% (3) 12.5% (2) 12.5% (1)
Table 7: Best Case Modification Distribution by Agents
Modification Type TheAIScientist AIDE Claude Code
Algorithmic Modification 75.0% (12) 83.3% (10) 37.5% (3)
Structural Modification 12.5% (2) 0.0% (0) 37.5% (3)
Data Processing 12.5% (2) 16.7% (2) 0.0% (0)
Parameter Configuration 0.0% (0) 0.0% (0) 25.0% (2)
Table 8: Best Case Modification Distribution by LLMs
Modification Type GPT-5 Gemini-2.5-pro Opus-4.1
Algorithmic Modification 14 (100.0%) 8 (57.1%) 3 (37.5%)
Structural Modification 0 (0.0%) 2 (14.3%) 3 (37.5%)
Data Processing 0 (0.0%) 4 (28.6%) 0 (0.0%)
Parameter Configuration 0 (0.0%) 0 (0.0%) 2 (25.0%)
Appendix C Protecting Evaluation Integrity
@@
Appendix D Additional Results
@@
This section provides a comprehensive comparative analysis of automated AI research systems, focusing on both their performance and operational efficiency. We evaluate TheAIScientist, AIDE, and Claude Code across eight core machine learning research problems, considering not only their final outcomes but also their operational behavior over time.
Additional analysis results covers multiple key dimensions: detailed exploration diversity (Tab. 9), operational reliability measured by step success rate (Tab. 10) and step completion rate (Tab. 11), computational efficiency in terms of total token usage (Tab. 12) and step-wise token usage (Tab. 13), runtime efficiency in terms of total time (Tab. 14) and step-wise time (Tab. 15). Together, these results provide detailed insights into the trade-offs between different agentic research designs, illustrating how they influence both the quality of research outcomes and the efficiency of resource utilization.
Table 9: Exploration Diversity Comparison of Best Performance Round. G2.5-Pro denotes Gemini-2.5-Pro. Bold numbers indicate the best (highest) performance in each row.
@@
Generalization 39.32 18.60 4.36 48.23 19.50
Data Eff. 28.18 20.19 18.03 10.92 10.08
Rep. Learn. 15.28 11.55 18.91 9.71 5.30
Cont. Learn. 18.86 24.83 36.72 18.80 2.71
Causality 24.03 21.37 38.61 11.65 6.43
Robust. & Rel. 45.44 24.04 43.03 21.23 16.19
Privacy 29.66 21.17 33.16 13.67 5.97
Fair. & Bias 26.88 25.14 35.97 13.10 3.84
Average $\pm$ Std 28.46±9.97 20.86±4.42 28.60±13.32 18.41±12.68 8.75±6.07
@@
Generalization 0.77 0.99 0.50 0.70 1.00
Data Eff. 0.96 0.65 0.23 0.29 0.94
Rep. Learn. 0.83 0.84 0.80 0.60 0.55
Cont. Learn. 0.96 0.94 0.82 0.92 1.00
Causality 0.98 0.89 0.88 0.98 1.00
Robust. & Rel. 0.98 0.89 0.87 0.79 1.00
Privacy 0.97 0.95 0.87 0.96 1.00
Fair. & Bias 0.18 0.21 0.18 0.04 0.13
Average $\pm$ Std 0.83±0.27 0.80±0.26 0.64±0.30 0.66±0.34 0.83±0.32
@@
Generalization 1.00 1.00 0.02 0.36 0.03
Data Eff. 1.00 1.00 0.30 1.00 0.16
Rep. Learn. 1.00 1.00 0.45 1.00 0.11
Cont. Learn. 1.00 1.00 0.28 0.70 0.02
Causality 1.00 1.00 1.00 0.52 0.11
Robust. & Rel. 1.00 1.00 1.00 0.56 0.01
Privacy 1.00 1.00 0.70 0.69 0.03
Fair. & Bias 1.00 1.00 0.39 0.7 0.08
Average $\pm$ Std 1.00±0.00 1.00±0.00 0.54±0.38 0.69±0.24 0.07±0.06
ML Problems TheAIScientist Claude Code
GPT-5 G2.5-Pro Opus-4.1
Generalization 9.34 9.88 5.98
Data Eff. 3.09 2.55 7.98
Rep. Learn. 4.69 3.99 8.89
Cont. Learn. 7.81 7.74 2.99
Causality 3.72 3.40 10.86
Robust. & Rel. 7.62 7.04 13.36
Privacy 6.64 4.45 13.88
Fair. & Bias 5.41 4.58 8.53
Average $\pm$ Std 6.04±2.18 5.45±2.51 9.06±3.64
@@
ML Problems TheAIScientist Claude Code
GPT-5 G2.5-Pro Opus-4.1
Generalization 0.09 0.10 1.99
Data Eff. 0.03 0.03 0.50
Rep. Learn. 0.05 0.04 0.81
Cont. Learn. 0.08 0.08 1.49
Causality 0.04 0.03 0.99
Robust. & Rel. 0.08 0.07 13.36
Privacy 0.07 0.04 4.63
Fair. & Bias 0.05 0.05 1.07
Average $\pm$ Std 0.06±0.02 0.05±0.03 3.10±4.34
@@
Fair. & Bias 5.20 2.87 5.32 3.60 0.30
Average $\pm$ Std 9.42±5.25 11.74±8.37 5.46±6.92 9.49±7.24 1.18±1.80
@@
Fair. & Bias 3.12 1.72 8.18 3.05 2.25
Average $\pm$ Std 11.05±10.71 14.23±14.55 12.55±12.17 10.48±8.29 13.26±11.29
Appendix E Prompts Details
This section presents the detailed prompt specifications that form the foundation of our automatic research agent framework. The prompts serve as the primary interface between human researchers and AI agents, translating high-level research objectives into actionable instructions that can guide systematic scientific inquiry across diverse machine learning domains. The prompt design philosophy centers on creating a structured yet flexible research environment that balances autonomy with scientific rigor. Rather than providing overly prescriptive instructions that limit creative exploration, these prompts establish clear boundaries, evaluation criteria, and operational constraints while encouraging the agent to develop and test novel hypotheses within established research paradigms.
@@ Appendix E Prompts Details
Second, we detail the Automatic Research Agent Framework that governs how agents interact with research codebases to conduct iterative experimentation. This operational framework transforms the conceptual research challenges into executable workflows, ensuring that agent behavior follows sound scientific methodology while maintaining reproducibility and experimental integrity. Together, these prompt specifications create a comprehensive research environment where automatic agents can contribute meaningfully to advancing machine learning across multiple disciplines, providing both the research contexts and the methodological framework necessary for systematic scientific progress.
@@
The following task descriptions establish comprehensive research contexts spanning 8 fundamental areas of machine learning. Each prompt follows a structured format that defines: 1) the researcher’s identity and expertise, 2) the specific technical setup including datasets and baseline methods, 3) clear optimization objectives and constraints, and 4) fairness criteria to ensure meaningful comparisons.
These prompts span a wide range of machine learning challenges (e.g., generalization, data efficiency, privacy, fairness, and robustness). Each task is anchored in established benchmarks and frameworks, such as DomainBed for domain generalization, EasyFSL for few-shot learning, and AIF360 for fairness-aware learning, providing realistic experimental environments that closely mirror actual research workflows.
The prompts are carefully crafted to encourage both incremental improvements to existing methods and the exploration of novel algorithmic approaches, while maintaining scientific rigor through controlled experimental conditions. This design enables systematic evaluation of how automatic agents can contribute to advancing the state-of-the-art across multiple ML disciplines simultaneously.
@@
System: You are an ambitious AI PhD student focused on advancing machine learning for causal inference, reasoning, and interpretable modeling. Task Description: You are working with the Dragonnet framework to estimate individual treatment effects (ITEs) in IHDP. The Infant Health and Development Program (IHDP) is a semi-synthetic benchmark dataset commonly used in causal inference, built from a real randomized trial with simulated treatment effects to enable ground-truth evaluation. Your goal is to improve the precision of treatment effect estimation across IHDP.
@@
System: You are an ambitious AI PhD student focused on improving fairness-aware learning with AIF360’s Adversarial Debiasing on the COMPAS dataset. Task Description: You are working with AIF360’s Adversarial Debiasing (classifier–adversary) as the baseline on the COMPAS dataset to evaluate the fairness–accuracy trade-off. Your goal is to minimize absolute Average Odds Difference toward parity (=0) while maintaining or improving Balanced Accuracy on held-out test splits and across protected subgroups (e.g., sex/race). You should enhance the baseline Adversarial Debiasing algorithm, but you may also propose entirely new fairness methods if they better support reduced absolute Average Odds Difference without sacrificing Balanced Accuracy. You are allowed to refine the classifier and adversary networks and the training pipeline, provided comparisons remain fair to the original setup (similar capacity, training budget, and data access). The priority is minimizing absolute Average Odds Difference while preserving or improving Balanced Accuracy.
Prompting Claude Code as an Automatic Research Agent